An ordinary (fair) die is a cube with the numbers 1 through 6 on the sides (represented by painted spots). Imagine that such a die is rolled twice in succession and that the face values of the two rolls are added together. This sum is recorded as the outcome of a single trial of a random experiment.
Compute the probability of each of the following events.
Event A: The sum is greater than 6.
Event B: The sum is an odd number.
(a.) P(A) =
(b.) P(B) =

Answer by mananth(16946)

.
Here the total possible outcomes = 36 n(S) =36
Event (A) Sum greater than 6

P(A) = n(A)/n(S) = 21/36 = 7/12
Event B: The sum is odd number
There are 18 possible outcomes n(B)=18
P(B)= n(B)/n(S)= 18/36 = 1/2
Convert to percent if required